Use the growth accounting equation to calculate productivity growth, given output growth of 4.5%, capital stock growth of 4%. labor employment growth of 3%, the output elasticity of capital of 0.3, and the output elasticity of labor of 0.7.
Answer #1

Answer

Growth accounting equation is given by :

%\Delta Y = %\Delta A + e_l%\DeltaY = %\DeltaA + eL(%\DeltaL) + eK(%\DeltaK)

where %\DeltaY = Output growth rate = 4.5%, %\DeltaK = capital stock growth of 4%, %\DeltaL = labor employment growth = 3%, eK = the output elasticity of capital = 0.3 and the eL = output elasticity of labor = 0.7. %\DeltaA = Productivity growth that we have to calculate.

Using above information and growth accounting equation we get :

4.5 = %\DeltaA + 0.3*4 + 0.7*3

=> %\DeltaA = 4.5 - (0.3*4 + 0.7*3) = 1.2

=> Productivity growth rate = %\DeltaA = 1.2%

Hence, Productivity growth = 1.2%